Creativeone24:

A system of equations is shown. -2a+2b = 24 3a+2b = 4 What is the value of a? Question 3 options: 4 28 -4 -5

dude:

First step: \(-2a+2b = 24\\ 3a+2b = 4\) Multiply the bottom equation by -1

Creativeone24:

ok thnk u is tht all??

dude:

No, I am only assisting you, this is not the answer

Creativeone24:

ok i did this step

dude:

Step 2: \(-2a+2b = 24\\ -3a-2b = -4\) Add both equations \[-5a=20\] Do you know what to do from now on?